OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form.

OpenStudy (anonymous):

3x – 3y = –8 x – 3y = –9 x – 3y = –8 3x + y = 9

OpenStudy (anonymous):

3x – 3y = –8 x – 3y = –9 3x + y = 9 x – 3y = –8

OpenStudy (anonymous):

it looks different

hartnn (hartnn):

the equation of line can be written as (y-y1)=m(x-x1) where (x1,y1) is a point on that line. so put any one of the points here and tell me what u get.

OpenStudy (anonymous):

(3, 1)

OpenStudy (anonymous):

@paulaaa

OpenStudy (anonymous):

@Hero can u save me pleeeeeeeeeease

hero (hero):

Plug each of the points one at a time into the equation y = mx + b

OpenStudy (anonymous):

how old r u i did that

hero (hero):

You should end up with 4 = 4m + b 3 = 1m + b

OpenStudy (anonymous):

i got that

hero (hero):

Now isolate b in each equation

hero (hero):

Let me know what you get

OpenStudy (anonymous):

i got -9

OpenStudy (anonymous):

sorry i didnt get notified intell know but it looks like u dont need my help u got hero so nvm

hero (hero):

You isolated b in each equation and got -9?

hero (hero):

You should have gotten b = 4 - 4m b = 3 - 1m

OpenStudy (anonymous):

i dont get it

OpenStudy (phi):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form. one way is find the slope: (4-3)/(4-1) = 1/3 find b (y intercept) y= mx + b with m= 1/3 use (1,3) , x=1, y=3 to find b 3= (1/3)*1+b b= 3- (1/3) = 8/3 so the equation is y= (1/3) x + 8/3 now put into standard form (x is first, positive coefficient)

hero (hero):

Plug each (x,y) point into the equation y = mx + b to get two equations: 4 = 4m + b 3 = 1m + b Isolate b in each equation to get b = 4 - 4m b = 3 - 1m set b = b to get 4 - 4m = 3 - 1m Place like terms on the same side to get 4 - 3 = 4m - 1m Simplify to get 1 = 3m Isolate m = to get 1/3 = m Subsitute m back into one of the equations above to solve for b: b = 3 - 1m b = 3 - 1(1/3) b = 9/3 - 1/3 b = 8/3 Now that you've found m and b, write the equation of the line in the form y = mx+ b: y = x/3 + 8/3